OpenStudy (anonymous):

if a,b,c are are in Arithmetic progression, prove bc-a^2, ca-b^2, ab-c^2 are also in AP.

terenzreignz (terenzreignz):

Well, if a, b, and c are in arithmetic progression, what immediately follows?

OpenStudy (anonymous):

b-a=c-b i did (ca-b^2)-( bc-a^2)=(ab-c^2)-(ca-b^2) but i am getting (a-b)= (b-c)

terenzreignz (terenzreignz):

And what's wrong with that? :P Does it not follow that if p = q then -p = -q ? ;)

OpenStudy (anonymous):

so i can multiply both sides with -1 ??

terenzreignz (terenzreignz):

Naturally :P 1 , 2, and 3 are in arithmetic progression... so 2 - 1 = 3 - 2 true enough, but also... 1 - 2 = 2 - 3 LOL ^_^

OpenStudy (anonymous):

thanks